OpenStudy (anonymous):

What is the solution to the equation 1/3 (9x-6) = 2x+3 ? x=?

OpenStudy (error1603):

1 Simplify — 3

OpenStudy (error1603):

Simplify 1/3

OpenStudy (anonymous):

how do u do that

OpenStudy (anonymous):

isnt already simplified

OpenStudy (error1603):

(1/3• (9x - 6)) - (2x + 3) = 0

OpenStudy (anonymous):

it eqauls 0

OpenStudy (error1603):

no

OpenStudy (anonymous):

oh im confused

OpenStudy (error1603):

3.1 Pull out like factors: 9x - 6 = 3 • (3x - 2)

OpenStudy (anonymous):

9x-3x

OpenStudy (error1603):

It's 2x+3

OpenStudy (error1603):

So it's (3x - 2) - (2x + 3) = 0

OpenStudy (anonymous):

3x-2x is x- -1 which turns into x+1 id understand i'm so lost i need help

OpenStudy (error1603):

You should have got x - 5 = 0

OpenStudy (error1603):

I'll just give it to you it's x = 5

OpenStudy (anonymous):

ok thanks

OpenStudy (error1603):

Any time